A compound has the composition 62.1% C, 10.3% H and 27.6% O. What is its empirical formula? (Ar: C = 12, H = 1, O = 16)


1
Expert's answer
2021-03-04T06:33:13-0500

Mass of C = 62.1 g

Moles of C = "\dfrac{62.1}{12}=5.175\space mol"

Mass of H = 10.3 g

Moles of H = 10.3 mol

Mass of O = 27.6 g

Moles of O = "\dfrac{27.6}{16}=1.725\space mol"

C = "\dfrac{5.175}{10.3}=0.50"

H = "\dfrac{10.3}{10.3}=1"

O = "\dfrac{1.725}{10.3}=0.167"

Multiplying by suitable integer 6,

"C=0.50\times 6=3\\H=1\times6=6\\O=0.167\times6=1.002\approx1"

Empirical formula = "C_3H_6O"